OpenStudy (anonymous):

EXPLAIN STEP BY STEP HOW U WULD FIND THE EXACT VALUE OF 3 OVER SQRT 8

OpenStudy (anonymous):



OpenStudy (ghazi):

\[\frac{ 3 }{ \sqrt 8 }= \frac{ 3 }{ \sqrt {4*2} }=\frac{ 3 }{ 2 \sqrt 2 }= \frac{ 3 }{ 2 \sqrt 2 }* \frac{ \sqrt 2 }{ \sqrt 2 }= \frac{3 \sqrt 2 }{ 4 }\]

OpenStudy (anonymous):

I WANT EXPLANATION IN WORDS LMAO

OpenStudy (ghazi):

@zahabaha clear??

OpenStudy (chihiroasleaf):

to find the value, first you should simplify the radicals, then rationalize the denominator..

OpenStudy (ghazi):

@nincompoop lol

OpenStudy (ghazi):

@zahabaha clear??

OpenStudy (anonymous):

ok well i hope i get the explanation right

OpenStudy (precal):

 in mathematics this is known as the exact value